The ePaint Store stocks paint in its warehouse and sells it online on its Internet Web site. The store stocks several brands of paint; however, its biggest seller is Sharman-Wilson Ironcoat paint, which has an estimated annual demand of 10,000 gallons of paint, an annual carrying cost of $0.75 per gallon, and an ordering cost of $150 per order. The company processes orders every day except on Sundays (i.e. 365-52 = 311 days).

1. Determine the optimal order size for Ironcoat paint. Enter number of gallons.

2. Compute the total annual cost for Ironcoat paint.

3. Calculate the number of orders that will be made annually.
Answer #1

Annual demand(D) = 10000 gallons

Annual carrying cost per gallon(H) = $0.75

Ordering cost (S) = $150

1) optimal order size(Q) = √(2DS/H)

= √[(2 X 10000 X 150) / 0.75]

= √(3000000/0.75)

= √4000000

= 2000 gallons

2) Annual carrying cost = (Q/2) H = (2000/2) 0.75 = (1000 X 0.75) = $750

Annual ordering cost = (D/Q) S = (10000/2000) 150 = (5 X 150) = $750

Total annual cost = Annual carrying cost + Annual ordering cost = $750 + $750 = $1500

3) Number of orders made annually = D/Q = 10000/2000 = 5 times